# Self-Improving Robust Preference Optimization

### v2

@@ 1 Introduction
(Step 2) Robust Preference Optimization of Generative Model: The next step is to exploit the self-improvement policy learned in the previous step to learn a generative LLM, $\pi$.
The key idea here is that the best generative policy can be identified as a policy that generates completions requiring minimal improvement using the optimal self-improvement policy $\pi_{\dagger}$ derived in step $1$. This goal can be achieved by minimizing the objective of step $1$ with respect to the generative policy for in-context completions, $y$. Similar to step $1$, this problem, in its KL-regularized form, can also be solved analytically in terms of the optimal improvement policy $\pi_{\dagger}$ and the optimal generative policy $\pi$. More significantly, we show that the solution for steps $1$ and $2$ can be estimated jointly through a single supervised direct preference optimization scheme using only a dataset of annotated pair-wise completions. Thus, one can solve both for the self-improvement policy $\pi_{\dagger}$ and $\pi$ by minimizing the supervised learning objective of SRPO. Unlike existing RLHF methods, this solution is independent of the behavior policy and is therefore robust to its changes.
@@ 4 Offline Solution for Optimizing SRPO Objective
$J^{*}(x)=\min_{\pi}\mathop{\mathbb{E}}_{xy_{1}\sim\pi(.|x)}\left[\beta(\log(Z^{*}(y_{1},x))+D_{\text{KL}}(\pi||\pi_{\text{ref}}|x))\right].$ (3)
@@ 4.1 Optimizing the Self-Improvement Policy $\pi_{\dagger}$
$+\mathop{\mathbb{E}}_{\begin{subarray}{c}(y_{l},y_{w},x)\sim\mathcal{D}\end{subarray}}\left[\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y_{w}|y_{w},x)}{\pi_{\text{ref}}(y_{w}|y_{w},x))}\right)-\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)\right]\right]^{2}$ (9)
@@ 4.2 Optimizing the Robust Generative Policy $\pi$
$\beta\log(Z^{*}(y_{1},x))=\beta(\log(\pi_{\text{ref}}(y_{1}|y_{1},x))-\log(\pi^{*}_{\dagger}(y_{1}|y_{1},x)))-\frac{1}{2}.$ (10)
@@
$\log(\pi^{*}_{\dagger}(y|x))=\log\left(\frac{\pi_{\text{ref}(y|x)}}{\pi_{\text{ref}}(y|y,x)}\pi^{*}(y|y,x)\right)-\log(Z^{*}(x)).$ (13)
@@
$\frac{p(y_{2}\succ y_{1}|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y_{2}|y_{1},x)}{\pi_{\text{ref}}(y_{2}|y_{1},x)}\right)$ $=\log\left(\frac{\pi_{\text{ref}}(y_{1}|x))}{\pi^{*}(y_{1}|x)}\right)-\log(Z^{*}(x)),$ (17)
$\frac{p(y_{1}\succ y_{2}|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)$ $=\log\left(\frac{\pi_{\text{ref}}(y_{2}|x))}{\pi^{*}(y_{2}|x)}\right)-\log(Z^{*}(x)).$ (18)
@@
$-\left(\log\left(\frac{\pi^{*}_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)-\log\left(\frac{\pi^{*}(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x))}\right)\right)\Bigg{]}.$ (20)
@@
$-\left(\log\left(\frac{\pi_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)-\log\left(\frac{\pi(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x))}\right)\right)\bigg{]}\Bigg{]}^{2}.$ (22)
@@
$-\left(\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)+\log\left(\frac{\pi(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x))}\right)\right)\bigg{]}-1\Bigg{]}^{2}.$ (24)
@@ 5 Robustness of SRPO
$P=\begin{pmatrix}0.5&0.99&0.3\\ 0.01&0.5&0.25\\ 0.7&0.75&0.5\end{pmatrix},$ (27)
@@ 6 Related works
Offline preference optimization was introduce as an alternative to more classic RLHF approaches, such as PPO (Schulman et al., 2017; Ouyang et al., 2022) or more generally policy-gradient-based approaches (Roit et al., 2023; Ahmadian et al., 2024). These methods require training a reward model on a preference dataset, usually with a Bradley-Terry model (Bradley & Terry, 1952). The reward model is then used to fine-tune the LLM via online RL, requiring many generations from the model. This reward model shares the common issue of DPO and other direct preference alignment methods, it is dependent on the sampling distribution $\mu$ used for constructing the preference dataset, contrary to SRPO. Moreover, classic RLHF is online, while SRPO is offline and thus more easily scalable.

### v3

@@ 1 Introduction
(Step 2) Robust Preference Optimization of Generative Model: The next step is to exploit the self-improvement policy learned in the previous step to learn a generative LLM, $\pi$. The key idea here is that the best generative policy can be identified as a policy that generates completions requiring minimal improvement using the optimal self-improvement policy $\pi_{\dagger}$ derived in step $1$. This goal can be achieved by minimizing the objective of step $1$ with respect to the generative policy for in-context completions, $y$. Similar to step $1$, this problem, in its KL-regularized form, can also be solved analytically in terms of the optimal improvement policy $\pi_{\dagger}$ and the optimal generative policy $\pi$. More significantly, we show that the solution for steps $1$ and $2$ can be estimated jointly through a single supervised direct preference optimization scheme using only a dataset of annotated pair-wise completions. Thus, one can solve both for the self-improvement policy $\pi_{\dagger}$ and $\pi$ by minimizing the supervised learning objective of SRPO. Unlike existing RLHF methods, this solution is independent of the behavior policy and is therefore robust to its changes.
@@ 4 Offline Solution for Optimizing SRPO Objective
$J^{*}(x)=\min_{\pi}\mathop{\mathbb{E}}_{y_{1}\sim\pi(.|x)}\left[\beta(\log(Z^{*}(y_{1},x))+D_{\text{KL}}(\pi||\pi_{\text{ref}}|x))\right].$ (3)
@@ 4.1 Optimizing the Self-Improvement Policy $\pi_{\dagger}$
$+\mathop{\mathbb{E}}_{\begin{subarray}{c}(y_{l},y_{w},x)\sim\mathcal{D}\end{subarray}}\left[\frac{1}{2}-\beta\left[\log\left(\frac{\pi_{\dagger}(y_{w}|y_{w},x)}{\pi_{\text{ref}}(y_{w}|y_{w},x))}\right)-\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)\right]\right]^{2}$
@@ 4.2 Optimizing the Robust Generative Policy $\pi$
$\beta\log(Z^{*}(y_{1},x))=\beta(\log(\pi_{\text{ref}}(y_{1}|y_{1},x))-\log(\pi^{*}_{\dagger}(y_{1}|y_{1},x)))-\frac{1}{2}.$
@@
$\log(\pi^{*}(y|x))=\log\left(\frac{\pi_{\text{ref}(y|x)}}{\pi_{\text{ref}}(y|y,x)}\pi^{*}_{\dagger}(y|y,x)\right)-\log(Z^{*}(x)).$
@@
$\frac{p(y_{2}\succ y_{1}|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y_{2}|y_{1},x)}{\pi_{\text{ref}}(y_{2}|y_{1},x)}\right)$ $=\log\left(\frac{\pi_{\text{ref}}(y_{1}|x))}{\pi^{*}(y_{1}|x)}\right)-\log(Z^{*}(x)),$
$\frac{p(y_{1}\succ y_{2}|x)}{\beta}-\log\left(\frac{\pi^{*}_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)$ $=\log\left(\frac{\pi_{\text{ref}}(y_{2}|x))}{\pi^{*}(y_{2}|x)}\right)-\log(Z^{*}(x)).$
@@
$-\left(\log\left(\frac{\pi^{*}_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)-\log\left(\frac{\pi^{*}(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x))}\right)\right)\Bigg{]}.$
@@
$-\left(\log\left(\frac{\pi_{\dagger}(y_{1}|y_{2},x)}{\pi_{\text{ref}}(y_{1}|y_{2},x)}\right)-\log\left(\frac{\pi(y_{2}|x)}{\pi_{\text{ref}}(y_{2}|x))}\right)\right)\bigg{]}\Bigg{]}^{2}.$
@@
$-\left(\log\left(\frac{\pi_{\dagger}(y_{l}|y_{w},x)}{\pi_{\text{ref}}(y_{l}|y_{w},x)}\right)+\log\left(\frac{\pi(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x))}\right)\right)\bigg{]}-1\Bigg{]}^{2}.$
@@ 5 Robustness of SRPO
$P=\begin{pmatrix}0.5&0.99&0.3\\ 0.01&0.5&0.25\\ 0.7&0.75&0.5\end{pmatrix},$
@@ 6 Related works
Offline preference optimization was introduced as an alternative to more classic RLHF approaches, such as PPO (Schulman et al., 2017; Ouyang et al., 2022) or more generally policy-gradient-based approaches (Roit et al., 2023; Ahmadian et al., 2024). These methods require training a reward model on a preference dataset, usually with a Bradley-Terry model (Bradley & Terry, 1952). The reward model is then used to fine-tune the LLM via online RL, requiring many generations from the model. This reward model shares the common issue of DPO and other direct preference alignment methods, it is dependent on the sampling distribution $\mu$ used for constructing the preference dataset, contrary to SRPO. Moreover, classic RLHF is online, while SRPO is offline and thus more easily scalable.